OpenStudy (anonymous):

In this discussion assignment, you will create an inverse variation problem. Then you will solve a problem that another student created. Before beginning this assignment, read over the problems you’ve studied so far in the Rational Expressions unit. Then make up a problem similar to the inverse variation problems. Make sure that you know how to solve the problem you create. Other students will post responses to your problem and you must be able to answer the other students. Review the problems that other students have posted. Select one and solve it. After you’ve completed these steps, post a

OpenStudy (anonymous):

Y=6 when x=3.find y when x=9???

OpenStudy (anonymous):

6-3= 3 this shows that y is 3 times greater than x that means 3y= x if x= 9 then 3y=9 y=9/3 y=3 i think this is how u'll do this

OpenStudy (anonymous):

i need a new one